OpenStudy (anonymous):

Linear Equation: A printer will produce 80 wedding invitations for $210. The price to produce 120 invitations is $290. The printer uses a linear function to determine the price for producing different amounts of invitations. How much will the printer charge to produce 60 invitations?

OpenStudy (campbell_st):

well you have 2 data points (80, 210) and (120, 290) find the slope of the line... to start then use y = mx + b substitute the slope.... then substitute either point to find b then you'll have the equation of the line. lastly, substitute x = 60 to find y.... or the cost

OpenStudy (anonymous):

I got: y=2(60)+80=200

OpenStudy (campbell_st):

well I think the slope is right but the value of b is wrong y = 2x + b so substitute x = 80 and y = 210 so 210 = 2(80) + b so b = 50 then the equation is y = 2x + 50 now substitute x = 60...